OpenStudy (anonymous):

solve for the y -4y+5/3 = -4/5y-7/5

OpenStudy (anonymous):

\[-4y+5/3 = -4/5y-7/5\] \[5/3 +7/5=-4/5y +4y\] \[46/15=-4/5y +4y\] \[230y/15=-4+20y \] \[230y =-60+300y\] \[60=300y-230y\] \[60=70y\] y=6/7

OpenStudy (anonymous):

Is it 4/5 * y or 4 over 5y?

OpenStudy (anonymous):

the forth step 230y/15=−4+20y ^2 you miss ^2

OpenStudy (anonymous):

60=300y^2−230y compleet from her you 'll get tow solutions